OpenStudy (anonymous):

(sec3x-cos3x)/(secx-cosx)=sec2x+1+cos2x

OpenStudy (anonymous):

sec2x - cos2x = sec2x + 1 + cos2x -cos2x = 1 + cos2x -1 = 2cos2x cos x = -1/4 x = acos -1/4 x = 1.8235